### v1

# A route to quantum computing through the theory of quantum graphs

Abstract
Based on our previous works, and in order to relate them with the theory of quantum graphs and the quantum computing principles, we once again try to introduce some newly developed technical structures just by relying on our toy example, i.e. the coordinate ring of $n\times n$ quantum matrix algebra $M_{q}(n)$, and the associated directed locally finite graphs $\mathcal{G}(\Pi_{n})$, and the Cuntz-Krieger $C^{*}$-graph algebras. Meaningly, we introduce a $(4i-6)$-quantum entangled system by using the Cuntz-Krieger $\mathcal{G}(\Pi_{i})$-families associated to the $64i-$ 6 distinct Hamiltonian paths of $\mathcal{G}(\Pi_{i})$, for $i\in\{2,\cdots,n\}$.
MSC Numbers (2020): Primary 46L05, 68Q09; Secondary 46L67, 81P40, 46L55, 17B81, 81P45.
Keywords: quantum computing; quantum information theory; quantum state; quantum system; qubit; quantum graph; Cuntz-Krieger $C^{*}$-graph algebra; quantum permutation group; Hamiltonian path.
1 An invitation to quantum graphs
Non-commutative Geometry is all about finding quantum (i.e. non-commutative) generalizations of some familiar structures, such as quantum topology, quantum probability (free probability), quantum groups, quantum permutation groups, and quantum information theory, etc., which triggers the following question:
Question 1.1.
What is the quantum analogue of a graph?
This problem has been studied from many perspectives, but the notion of quantum graphs (also called non-commutative graphs in [9]) first has been coined and introduced by Duan, Severini, Winter [9] in terms of operator systems as the confusability graph of a quantum channel in quantum information theory. As an analogue of the fact that simple undirected classical graphs are irreflexive symmetric relations, Weaver [10] formulated quantum graphs as reflexive symmetric quantum relations on a von Neumann algebra, which somehow extends the results obtained in [9], and following that quantum relations were introduced by Kuperberg, Weaver [4]. Later on, following these works, Musto, Reutter, Verdon [5] formulated finite quantum graphs as the adjacency operators on tracial finite quantum sets, and Brannan et al. In [2] a generalization has been introduced for non-tracial settings.
Set $\Gamma=(\Gamma^{0},\Gamma^{1})$ be a (directed) graph with vertex set $\Gamma^{0}=\{1,\cdots,n\}$, and the edge set $\Gamma^{1}=\{e_{ij}=(i,j)\mid\ \text{if}\ i\sim j\}$, where $\sim$ has the usual meaning that the vertices $i$ and $j$ are connected by edge $e_{ij}$. Then there are three different approaches (as we know) to the theory of quantum graphs, and they could be indicated as follows:
  1. I)
    The First approach will be applied by quantizing the confusability graph of the classical channels. This could be done by using the matrix quantum graphs and the operator systems, by projecting $P_{\mathfrak{S}}$ onto the operator system $\mathfrak{S}$.[9]
  2. II)
    The second approach will be applied by quantizing the edge set $\Gamma^{1}$, by using the quantum relations and the projection $P_{\Gamma^{1}}$ from $\chi_{\Gamma^{1}}$.[10]
  3. III)
    The third approach will be applied by quantizing the adjacency matrix, by using the categorical theory of the quantum sets and the quantum functions, and projecting $P_{\Gamma}$ by using the adjacency matrix $A_{\Gamma}$.[5]
It has been proved that under some appropriate identifications, the range of the above projections will be the same operator system!
In this paper, we will apply the first approach (I) to the quantum graphs by using the quantum confusability graph associated with the quantum channel applicable to our initial example $\mathcal{G}(\Pi_{2})$ based on our toy example.
2 Notation
Throughout this paper, $\mathbb{K}$ will stand for the ground field, and it will be assumed to be arbitrary unless otherwise stated.
We often will simply denote the matrix algebra $M_{d}(\mathbb{C})$ of $d\times d$ complex-valued matrices, with $M_{d}$, and the identity $d\times d$ matrix with $I_{d}$, and $M_{q}(n)$ will stand for the space of $n\times n$ quantum matrices, and by $\mathbb{K}[M_{q}(n)]$ we mean the ring of coordinate ring of $M_{q}(n)$.
$\mathcal{G}_{n}$ will stand for the directed locally finite connected graph associated with the defining relations of $\mathbb{K}[M_{q}(n)]$, and $\Pi_{n}$ represents the adjacency matrix of $\mathcal{G}_{n}$. After that, we will consider $\mathcal{G}_{n}:=\mathcal{G}(\Pi_{n})$.
By $u=(u_{ij})_{i,j}$, we mean the magic unitary matrix and $\pi_{n}$ will stand for the unique commuting magic unitary matrix with $\Pi_{n}$.
For a Hilbert space $\mathcal{H}$, we will consider the set of all (bounded) linear operators from $\mathcal{H}$ to $\mathbb{K}$, with $B(\mathcal{H},\mathbb{K})$, and we write $B(\mathcal{H})\equiv B(\mathcal{H},\mathcal{H})$ for the set of all linear operators acting on $\mathcal{H}$.
We use * to present the complex conjugate transpose of vectors and matrices.
In some places, we will use the Dirac notation from quantum mechanics, meaning that a column vector which is called a “ket” and is written as $|\psi\rangle\in\mathbb{C}^{d}$, and $\langle\psi|\equiv|\psi\rangle^{*}$ to indicate its dual.
By $\{0,1\}^{q}$, we mean the binary strings of length $q$.
3 Quantum Graphs, $C^{*}$-algebraic approach
Even though the various approaches to quantum graphs, mentioned at the end of section 1, look quite different (meaning that they are from different perspectives. For instance they are from the $C^{*}$-algebraic perspective, operator spaces, and from the Frobenius algebras point of view), it is known that they ultimately will lead to the same non-commutative theory.
In this paper, we will focus on a $C^{*}$-algebraic description of the theory of quantum graphs based on [5].
To start, let $\Gamma=(\Gamma^{0},\Gamma^{1})$ be as before, for $\Gamma^{0}$ the set of vertices and $\Gamma^{1}$ the set of edges. We have the following definition, quoted from [9].
Definition 3.1.
The quantum (non-commutative) graph associated with $\Gamma$ will be defined as the operator system
$\mathfrak{S}_{\Gamma}:=\,\mathsf{span}\{E_{ij}:(i,j)\in\Gamma^{1}\ \text{or}\ i=j,\forall i,j\in\Gamma^{0}\}\subseteq M_{n},$
for $E_{ij}$ the matrix units in $M_{n}$.
More generally we have the following definition.
Definition 3.2.
[5] In general, any such operator system in $M_{n}$ will be called a matrix quantum graph.
Next, we need another important concept which is called a quantum channel, and it is a completely positive and trace-preserving linear map defined as follows.
Definition 3.3.
[9] In quantum information theory, a quantum channel is represented by a completely positive trace-preserving map $\Psi:M_{i}\to M_{j}$, written in the Choi-Kraus form
$\Psi(\mathcal{X})=\sum_{m}S_{m}\mathcal{X}S_{m}^{*}\qquad\ \text{with}\ \qquad\sum_{m}S_{m}^{*}S_{m}=I_{i},$ (3.4)
for every $\mathcal{X}\in M_{i}$, where $I_{i}$ is the identity matrix in $M_{i}$.
As has been pointed out in section 1, there are three known approaches to the theory of quantum graphs (as we know), and the one which we are going to follow, is related to the confusability graph, and of course, the noncommutative version. Hopefully, here we are not required to quantize the classical version, as ours is already quantized!
Let us quote the following theorem from [9].
Theorem 3.5.
For a quantum channel $\Psi:M_{n}\to M_{m}$, and any $\mathcal{X}\in M_{n}$, with $\Psi(\mathcal{X})=\sum_{i=1}^{r}S_{i}\mathcal{X}S_{i}^{*}$, the confusability graph of $\Psi$ is the operator system
$\mathcal{S}_{\Psi}=\,\mathsf{span}\{S_{i}^{*}S_{i}:1\leq i,j\leq r\}\subseteq M_{n}.$ (3.6)
It is known that quantum channels generalize the unitary evolution of the isolated quantum systems to the open quantum systems, on which by looking at them from the mathematical point of view, a quantum channel can be described as a linear completely positive trace-preserving map $\Psi$ from $M_{d}$ to itself, and in this paper, by a quantum channel we mean these kinds of maps!
Note that the trace preservation condition is almost essential, since quantum channels are usually meant to map density matrices to the density matrices, and the complete positivity condition could be stated in a way that for any $m\geq 1$, $\Psi\otimes I_{d}:M_{dm}(\mathbb{C})\to M_{dm}(\mathbb{C})$ is a positive map.
The following equivalent three characterizations of the quantum channel $\Psi$ from $M_{d}$ to $M_{d}$ are known and have been quoted from [9]:
  1. i)
    Stinespring dilation. For every $\mathcal{X}\in M_{m}$, there exists a finite dimensional Hilbert space $\mathcal{H}$, of dimension $n=m^{2}$, and an isometry $S\in M_{m}$ acting on $\mathcal{H}$, i.e. $S:\mathbb{C}^{m}\to\mathbb{C}^{m}\otimes\mathbb{C}^{n}$, such that
    $\Psi(\mathcal{X})=\left(I_{m}\otimes Tr_{n}\right)(S\mathcal{X}S^{*}).$ (3.7)
  2. ii)
    Choi-Kraus decomposition. For every $\mathcal{X}\in M_{m}$, there exists an integer $k$ and isometries $S_{1},\cdots,S_{k}\in M_{m}$ satisfying $\sum_{i}S_{i}^{*}S_{i}=I_{m}$, such that
    $\Psi(\mathcal{X})=\sum_{i=1}^{n}S_{i}\mathcal{X}S_{i}^{*}.$ (3.8)
  3. iii)
    Choi matrix. Set $n=m^{2}$. The positive semidefinite Matrix $C_{\Psi}\in M_{n}$, satisfying $\left(I_{m}\otimes Tr_{n}\right)(C_{\Psi})=I_{m}$, is called the Choi matrix of $\Psi$ such that
    $C_{\Psi}:=\sum_{i,j=1}^{d}E_{ij}\otimes\Psi(E_{ij})\in M_{m}\otimes M_{m}.$ (3.9)
In this paper, we will follow the second characteristic (ii), which has been already proven that is equivalent to other characteristics (i), (iii).[9]
4 $C^{*}$-Graph Algebras, a rout to quantum ($C^{*}$-algebraic) graphs
By following the constructions in [7], and considering the finite locally connected (directed) graph $\mathcal{G}_{2}:=\mathcal{G}(\Pi_{2})=(\mathcal{G}_{2}^{0},\mathcal{G}_{2}^{1})$, associated with the coordinate ring of the $2\times 2$-quantum matrix algebra $M_{q}(2)$, let us consider its set of vertices and edges, exactly as in [7], such that $\mathcal{G}_{2}^{0}=\{x_{11}:=u,x_{12}:=v,x_{22}:=k,x_{21}:=w\}$ and $\mathcal{G}_{1}^{1}=\{x_{11}\overrightarrow{\sim}x_{12}:=e,x_{11}\overrightarrow{\sim}x_{21}:=f,x_{12}\overrightarrow{\sim}x_{22}:=h,x_{21}\overrightarrow{\sim}x_{22}:=g,x_{12}\overrightarrow{\sim}x_{21}:=i,x_{21}\overrightarrow{\sim}x_{12}:=j\}$, illustrated as in figure 1.
$u$$v$$k$$w$ehfgij
Figure 1: Directed locally connected graph related to $\Pi_{2}$
In [7], we proved that there is an infinite-dimensional $C^{*}$-graph algebra associated with $\mathcal{G}(\Pi_{2})$ as follows.
Proposition 4.1.
By considering the underlying infinite dimensional Hilbert space $\mathcal{H}:=\ell^{2}(\mathbb{N})$, the set
[…]
$S_{g}:=\sum_{n=1}^{\infty}E_{6n-4,3n-1},S_{i}:=\sum_{n=1}^{\infty}E_{6n-1,3n},S_{j}:=\sum_{n=1}^{\infty}E_{6n-3,3n-1}\}$ (4.2)
will define a Cuntz-Krieger $\mathcal{G}(\Pi_{2})$-family and will give us an infinite dimensional graph $C^{*}$-algebra structure $\mathcal{C}^{*}(\Pi_{2})$.
Later on, we will use the $\mathcal{G}(\Pi_{2})$-Cuntz-Krieger family $S$, introduced in Proposition 4.1, in order to define the associated quantum channel $\Psi_{\mathcal{G}(\Pi_{2})}$ and the noncommutative confusability graph $\mathcal{S}_{\Psi_{\mathcal{G}(\Pi_{2})}}$, and finally we will propose the associated 2-qubit quantum system, and the higher orders.
4.1 A route to quantum computing through the theory of quantum graphs
The field of quantum error correction is concerned with protecting fragile quantum information from the unexpected errors which might happen while processing large-scale computations, in an effort to build a much more efficient quantum computer than their predecessors. In February 2019, Adrian Chapman while looking for some new approaches in order to tackle a long-standing search for the Holy Grail of quantum error correction, noticed a critical relation between the theory of quantum graphs (from the $C^{*}$-algebraic point of view) and the previous findings concerning the quantum error corrections. He tried to develop a new way of encoding quantum information that is resistant to errors by constructions and doesn’t require active correction. What we do, is also somehow the same on an epsilon scale of what has been done, on which might be helpful, and could be simply stated as follows!
Following the above-stated statements, and in order to relate the core base of our work with the theory of quantum computation and the error corrections, let us invite $\mathcal{G}(\Pi_{2})$ back to the scene, and consider its set of Cuntz-Krieger family (4.2).
Consider figure 1, and note that $\mathcal{G}(\Pi_{2})$ consists of two distinct Hamiltonian paths. Let us call them $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$, such that $\mathcal{P}_{1}$ starts from the vertex $u$ and passes through the edges $e,i,g$ and terminates in the vertex $k$. On the other hand, $\mathcal{P}_{2}$ starts from the vertex $u$ and passes through the edges $f,j,h$ and terminates in the vertex $k$, as has been demonstrated.
Then by employing the defining relations (4.2), it will not be too difficult to see that $S_{e}^{*}S_{e}+S_{i}^{*}S_{i}+S_{g}^{*}S_{g}=I$, (and $S_{f}^{*}S_{f}+S_{j}^{*}S_{j}+S_{h}^{*}S_{h}=I$) are true statements within $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$ respectively, and then by using the definition 3.3, we might be interested in defining a completely positive trace-preserving map $\Psi_{\mathcal{G}(\Pi_{2})}^{1}:M_{m}\to M_{m}$ (and $\Psi_{\mathcal{G}(\Pi_{2})}^{2}:M_{m}\to M_{m}$); where $m$ will be specified later; in a way that we have $\Psi_{\mathcal{G}(\Pi_{2})}^{1}(\mathcal{X})=\sum_{m}S_{m}\mathcal{X}S_{m}^{*}$ (and $\Psi_{\mathcal{G}(\Pi_{2})}^{2}(\mathcal{X})=\sum_{m}S_{m}\mathcal{X}S_{m}^{*}$) for all $\mathcal{X}\in M_{m}$, and we will call it our quantum channel. After that, for $\ell\in\{e,g,i\}$ (and $\ell\in\{f,h,j\}$), the generating partial (matrix) isometries $S_{\ell}$, will be called the Choi-Kraus operators, and the noncommutative (confusability) graph of $\Psi_{\mathcal{G}(\Pi_{2})}^{1}$ (and $\Psi_{\mathcal{G}(\Pi_{2})}^{2}$) will be the operator system
$\mathcal{S}_{\Psi_{\mathcal{G}(\Pi_{2})}^{1}}=\,\mathsf{span}\{S_{\ell}^{*}S_{\ell}\}\subseteq M_{m},$ (4.3)
$\left(\text{and}\ \mathcal{S}_{\Psi_{\mathcal{G}(\Pi_{2})}^{2}}=\,\mathsf{span}\{S_{\ell}^{*}S_{\ell}\}\subseteq M_{m},\right)$ (4.4)
in a way that completely characterizes the number of zero-error messages on which one can send through the quantum channel $\Psi_{\Pi_{2}}^{1}(\text{and}\ \Psi_{\Pi_{2}}^{2})$.
Lemma 4.5.
There are two distinct quantum channels associated with $\mathcal{G}(\Pi_{2})$.
Proof.
As stated above, in $\mathcal{G}(\Pi_{2})$ there are two distinct Hamiltonian paths $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$, such that we have $S_{e}^{*}S_{e}+S_{i}^{*}S_{i}+S_{g}^{*}S_{g}=I$, and $S_{f}^{*}S_{f}+S_{j}^{*}S_{j}+S_{h}^{*}S_{h}=I$, respectively, and it is known that associated to each of these relations, there exists quantum channels $\Psi_{\Pi_{2}}^{1}$ and $\Psi_{\Pi_{2}}^{2})$, and the correspondent noncommutative (confusability) graphs $\mathcal{S}_{\Psi_{\mathcal{G}(\Pi_{2})}^{1}}$ and $\mathcal{S}_{\Psi_{\mathcal{G}(\Pi_{2})}^{2}}$. ∎
Following the above findings, and in order to move a little bit further, and somehow alternate the direction to the theory of quantum computing, we use the relation between the degree of sink and source nodes and the number of distinct Hamiltonian paths in our directed graphs $\mathcal{G}(\Pi_{n})$, in order to predict the accuracy of the possible quantum systems!
But before that, we note that the above statement, on the existence of a relation between the degree and the number of distinct Hamiltonian paths, is not always true! There are such relations, exactly when the source and the sink vertices have the same degree, and else there is no such relation.
Lemma 4.6.
The number of distinct Hamiltonian paths in $\mathcal{G}(\Pi_{n})$ is equal to $64n-$ 6.
Proof.
We will prove this assertion by combining the recursive relation on $\mathcal{D}_{n}$ and the induction processes on $n$.
To proceed with the proof, let us start by considering $\mathcal{G}(\Pi_{2})$. Note that in $\mathcal{G}(\Pi_{2})$ the degree is equal to $\mathcal{D}_{2}:=\mathcal{D}_{\mathcal{G}(\Pi_{2})}=2=2(2-1)$. And as has been stated before, there are exactly $2=2(2-1)=2(\mathcal{D}_{2}-1)$ distinct Hamiltonian paths, namely $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$.
However, in $G(\Pi_{3})$, the degree is $\mathcal{D}_{3}:=\mathcal{D}_{\mathcal{G}(\Pi_{3})}=4$, and once again, by doing some precaution, it is easy to see that $\mathcal{D}_{3}=2(3-1)$. And in accordance, it is easy to see that it possesses $6=2(4-1)=2(\mathcal{D}_{3}-1)$ distinct Hamiltonian paths. This algorithm continues until $G(\Pi_{n})$, but let us prove it. We have
$\mathcal{D}_{2}=2=2(2-1),$
$\mathcal{D}_{3}=4=2(3-1)=2(2-1)+2,$
$\mathcal{D}_{4}=6=2(4-1)=2(2-1)+2+2$
$\mathcal{D}_{5}=8=2(5-1)=2(2-1)+2+2+2,$
and hence, we obtain the recursive relation $\mathcal{D}_{i}=\mathcal{D}_{2}+\mathcal{D}_{i-1}$, for any $i\in\{2,\cdots,n\}$.
Suppose the first step, i.e. $i=2$, and the $n$th step, i.e. $\mathcal{D}_{n}$ satisfies. We have
$\mathcal{D}_{n+1}=\mathcal{D}_{2}+\mathcal{D}_{n}$ $=2(2-1)+2(n-1)$
$=2(2)-2+2(n)-2$
$=2n$
$=2(n+1-1).$
And since, by definition, we have the number of distinct Hamiltonian paths #$\#\mathcal{H}_{\mathcal{G}(\Pi_{n})}=2(\mathcal{D}_{n}-1)$, and hence we have #$\#\mathcal{H}_{\mathcal{G}(\Pi_{n})}=2(2(n-1)-1)=4n-6$.
And we have the following result.
Lemma 4.7.
There are $64n-$ 6 distinct quantum channels associated with $\mathcal{G}(\Pi_{n})$.
Proof.
The proof of this lemma, is almost clear, since by lemma 4.6, we have that the number of distinct Hamiltonian paths in $\mathcal{G}(\Pi_{n})$ is equal to $64n-$ 6, and we know that to each Hamiltonian path, we can associate a quantum channel and a correspondent noncommutative confusability graph. ∎
Then by employing these paths and the partial (matrix) isometries associated with each edge substituted in these Hamiltonian paths, we can formulate the associated quantum channel and the noncommutative (confusability) graph of $\mathcal{G}(\Pi_{n})$, and after that, by associating to the vertices involved in each Hamiltonian path we might get some results concerning the $64n-$ 6-qubit quantum system, which is the goal!
The idea of using (distinct) Hamiltonian paths, came from the famous travelling salesman problem (TSP) and its generalization, i.e. the travelling salesman path problem (TSPP), and trying to find an optimized algorithm in order to solve that. Since the matrix representation of $\mathbb{K}[M_{q}(n)]$ possesses a very nice triangulated partitioning, meaning that the associated directed graphs $\mathcal{G}(\Pi_{n})$ are triangulated, and as it is known there is a very common way in studying the TSP by using polygonometric (polyhedral) methods, and this has triggered us to propose the following way in approaching these kinds of problems!
4.1.1 2-qubit quantum system
Qubit is the basic unit in the quantum information theory, and in order to formally define it, we require the following first initial postulate of quantum mechanics.
Postulate 4.8 (State Space Postulate).
[1] In any isolated (physical) system, there is a complex space together with an inner product known as the state space of the system.
Remark 4.9.
Any physical system will completely be described by its unit vector from its state space, on which the vector will be called the state vector of the system.
Depending on the underlying (directed) graph of the quantum system, the state space will be different. But let us just consider $\mathbb{C}^{p}$ to be the state space of our pre-assumed quantum system. Note that, depending on the choice, the state vectors of the state space, will completely describe the space. By a state vector, we mean the unit vector in $\mathbb{C}^{p}$, which will be called a qubit, in the language of the information theory, and we have the following definition.
Definition 4.10.
A linear combination $\phi=c_{1}|0\rangle+\cdots+c_{p}|p-1\rangle$ of the basis vectors of $\mathbb{C}^{p}$, for $c_{1},\cdots,c_{p}\in\mathbb{C}$ will be called a qubit.
Remark 4.11.
In Definition 4.10, $c_{i}$s, for $i\in\{1,\cdots,p\}$, usually (in the literature) will be called amplitudes, and they satisfy in the condition $|c_{1}|^{2}+\cdots+|c_{p}|^{2}=1$, which is the usual normalization condition, it is not too difficult to see that the normalization condition will simply imply $\||\phi\rangle\|=\sqrt{\langle\phi|\phi\rangle}=1$.
At this point, we need the second postulate, which could be stated as follows and is concerned with qubit multiples.
Postulate 4.12.
[1] If a system $\phi$ is a combination of two different (physical) systems $\phi_{1}$ and $\phi_{2}$, with the corresponding states $\mathbb{E}_{1}$ and $\mathbb{E}_{2}$, then the state of $\phi$ will be $\mathbb{E}_{1}\otimes\mathbb{E}_{2}$, correspondent to $\phi_{1}\otimes\phi_{2}$.
Remark 4.13.
In the literature, mostly $\otimes$ symbol is omitted and $|\phi_{1}\rangle|\phi_{2}\rangle$ or $|\phi_{1}\phi_{2}\rangle$ is written in place of the composition of states $|\phi_{1}\rangle$ and $|\phi_{2}\rangle$. But, here in this paper we will not follow this notation, and we will refer to it as just the usual composition.
We have the following important definition, which has been quoted from [1].
Definition 4.14.
The state of a $q$-qubit quantum system is a unit vector in $(\mathbb{C}^{2})^{\otimes q}$.
Now, getting back to the main constructions, let $\{0,1\}^{q}$ be the binary strings of length $q$. Then by using the binary representation of the natural number $n=2^{k-1}a_{k-1}+\cdots+2a_{1}+a_{0}$, for $a_{j}\in\{0,1\}$, for all $10\leq j\leq k-$ 1, represented as $s=a_{k-1}a_{k-2}\cdots a_{1}a_{0}$, it is known that we can define a $q$-qubit system as [1]
$|\phi\rangle=\sum_{j\in\{0,1\}^{q}}\alpha_{j}|j\rangle,$
$\text{where}\ \sum_{j\in\{0,1\}^{q}}|\alpha_{j}|^{2}=1,$
for $\alpha_{j}\in\mathbb{C}$. Equivalently we have
$|\phi\rangle=\sum_{k=0}^{2^{q}-1}\alpha_{k}|k\rangle,$
$\text{where}\ \sum_{k=0}^{2^{q}-1}|\alpha_{k}|^{2}=1,$
for $\alpha_{k}\in\mathbb{C}$.
At this point, we already have what we need in order to enter quantum computing and quantum information theory!
So, let us invite $\mathcal{G}(\Pi_{2})$ back to the scene. Take $m=4$ in quantum channel $\Psi_{\mathcal{G}(\Pi_{2})}$. Here, our quantum system will be a combination of two subsystems $\mathbb{E}_{1}$ and $\mathbb{E}_{2}$ with the correspondent states $\phi_{1}$ and $\phi_{2}$, and since in this case, the Hamiltonian paths are a combination of three edges, hence let $\mathbb{C}^{3}$ be our state space $\phi$. Consider $\phi_{1}=\alpha_{1}|0\rangle+\alpha_{2}|1\rangle+\alpha_{3}|2\rangle$ and $\phi_{2}=\beta_{1}|0\rangle+\beta_{2}|1\rangle+\beta_{3}|2\rangle$ such that we have $\sum_{i=1}^{3}\alpha_{i}^{2}=\sum_{i=1}^{3}\beta_{i}^{2}=1$. Not that the only possible options for $\alpha_{i}$s and $\beta_{i}$s are ±$\pm\frac{1}{\sqrt{3}}$.
Now we can define a 2-qubit system associated with the pre-assumed quantum system as follows
$|\phi\rangle=\sum_{j\in\{0,1\}^{2}}\alpha_{j}|j\rangle=\alpha_{00}|00\rangle+\alpha_{01}|01\rangle+\alpha_{10}|10\rangle+\alpha_{11}|11\rangle$
$\hskip 14.22636pt=\alpha_{0}|0\rangle+\alpha_{1}|1\rangle+\alpha_{2}|2\rangle+\alpha_{3}|3\rangle,$
$\text{where}\ \sum_{j\in\{0,1\}^{2}}|\alpha_{j}|^{2}=1$
or, equivalently
$|\phi\rangle=\sum_{k=0}^{3}\alpha_{k}|k\rangle=\alpha_{0}|0\rangle+\alpha_{1}|1\rangle+\alpha_{2}|2\rangle+\alpha_{3}|3\rangle,$ (4.15)
$\text{where}\ \sum_{k=0}^{3}|\alpha_{k}|^{2}=1.$
It is easy to see that, in (4.15), the only possible options for $\alpha_{k}$s are ±$\pm\frac{1}{2}$, and as a result we have the following proposition.
Proposition 4.16.
Our pre-assumed 2-qubit quantum system associated with $\mathcal{G}(\Pi_{2})$ is entangled!
Proof.
The proof of this proposition is almost a trivial conclusion of the Definition 4.17, because, in any case, $|\phi\rangle$ will never be equal to the tensor product $|\phi_{1}\rangle\otimes|\phi_{2}\rangle$, because amplitudes in $|\phi_{1}\rangle\otimes|\phi_{2}\rangle$ in their best possible state will be equal to ±$\pm\frac{1}{3}$, while in $|\phi\rangle$ we have ±$\pm\frac{1}{2}$. ∎
Definition 4.17.
[1] A quantum state $|\phi\rangle\in(\mathbb{C}^{m})^{\otimes q}$ is a product state if it can be expressed as a tensor product $|\phi_{1}\rangle\otimes\cdots\otimes|\phi_{q}\rangle$ of $q$ 1-qubit states. Otherwise, it is entangled.
4.1.2 ($64n-$ 6)-qubit quantum system
We have the following claim.
Claim 4.18.
In the case of $\mathcal{G}(\Pi_{n})$, the pre-assumed $(4n-6)$-qubit quantum system is entangled.
Note that in order to prove or disprove this claim, the person first needs to prove or disprove Claim 2.7., from [7].
5 Concluding remarks
We believe that the research conducted in this paper is very interesting, and if we want to describe it in just one sentence it would be “from simplicity to complexity”!
Once again, and as in our previous works [6, 7], we started from our toy example $\mathbb{K}[M_{q}(n)]$, and motivated by the graph $C^{*}$-algebras through Cuntz-Krieger $\mathcal{G}(\Pi_{n})$-family for the finite locally connected directed graph $\mathcal{G}(\Pi_{n})$, and then we introduced the first initial concept of a 2-qubit entangled quantum system based on $\mathcal{G}(\Pi_{2})$, by discovering its relations with the Hamiltonian paths!
Then we proposed a claim on if the general case also provides us with an entangled quantum system.
In continuation, for an interested reader, a direction which could be proposed is to look at the other characteristics (i), (iii), and to prove that if they really are equivalent with (i), and since this is a structural example, based on construction, hence we believe that it is almost applicable!
6 Acknowledgement
The first author of this manuscript would like to express his sincere thanks and gratitude for the hospitality of the Department of Mathematics of the Azarbaijan Shahid Madani University, on which most of the work has been extracted and has been finalized, and he would like to acknowledge his postdoctoral grant with contract No. 117.d.22844 - 08.07.2023.
The research of the first author was also in part supported by a grant from IPM (No. 1403170014)
The second author was supported by the Department of Mathematics of the Azarbaijan Shahid Madani University under grant No. 1402/270 - 19.04.2023.
The authors of this manuscript have equal contributions to this work.